OpenStudy (anonymous):

Find all solutions of the equation 2sin2x−cosx=1 in the interval [0,2π), what is x? Find all solutions of the equation 2sin2x−cosx=1 in the interval [0,2π), what is x? @Mathematics

OpenStudy (anonymous):

hey its actually 2sin^(2)x- cosx=1

OpenStudy (jamesj):

That changes things. sin^2 x = 1 - cos^2 x So you can rewrite your equation as a quadratic equation in cos x.

OpenStudy (anonymous):

sin^2 x - cos^2x + 1?

OpenStudy (jamesj):

\[ 2\sin^2x−\cos x = 2(1 - \cos^2 x) + \cos x = -2\cos^2 x + \cos x + 2 \]

OpenStudy (jamesj):

Hence if \[ 2\sin^2 x−\cos x = 1 \] then \[ -2\cos^2 x + \cos x + 1 = 0 \]

OpenStudy (jamesj):

Now let \( u = \cos x \), then you solve the quadratic equation \[ 2u^2 - u - 1 = 0 \] for u and then find the x such that cos x equals those solutions.

OpenStudy (anonymous):

whoa, do you think you can demonstrate that. im not sure i know how to solve for u and x there

OpenStudy (jamesj):

You can solve \( 2u^2 - u - 1 = 0 \). Use the quadratic formula. Tell me what solutions you get.

OpenStudy (anonymous):

(1,-.5)

OpenStudy (anonymous):

x1= 1 x2= -.5

OpenStudy (jamesj):

Correct: \( u = 1, -1/2 \). Now, for what x is \[ \cos x = 1 \]or \[ \cos x = -1/2 \] where \( x \ \in \ [0,2\pi) \).

OpenStudy (anonymous):

is that saying x does not equal to 0,2pi

OpenStudy (jamesj):

Where x is in that interval. That's what you have as part of your original question; x is in that interval.

OpenStudy (anonymous):

so if i had to enter an answer for x, i would say what?

OpenStudy (jamesj):

"Find all solutions of the equation 2sin^2x−cosx=1 in the interval [0,2π)"

OpenStudy (jamesj):

For what x is cos x = 1? and for cos x = -1/2?

OpenStudy (anonymous):

this is exactly what my homework says Find all solutions of the equation 2sin2x−cosx=1 in the interval [0,2π). The answer is x= . Note: If there is more than one solution, enter them separated by commas. If needed enter π as pi.

OpenStudy (anonymous):

i dont know what they want me enter for x...should i say x is in that interval

OpenStudy (jamesj):

So far we have shown that cos x = 1 or cos x = -1/2. Now the question is for what x is that true. For example, for what values of x is it the case that cos x = 1?

OpenStudy (anonymous):

how do i find that?

OpenStudy (jamesj):

What is sin 0 ?

OpenStudy (anonymous):

0 right

OpenStudy (jamesj):

yes. What is sin(pi/2)?

OpenStudy (anonymous):

1

OpenStudy (jamesj):

Yes. For values of x is sin x = 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find all solutions of the equation sec^(2)x−2=0. The answer is A+Bkπ where k is any integer and 0<A<π/2, Find A and B can you help me find just B also pls James

OpenStudy (phi):

You might want to fine-tune your solution, as the equation you solve should be 2u^2 + u -1=0 , u =cos(x). I think you flipped a sign somewheres

OpenStudy (anonymous):

phi its is not accpeting the u in the equation

OpenStudy (jamesj):

Thanks Phi. Yes, if \[ 2\sin^2 x - \cos x = 1 \] then \[ 2 (1 - \cos^2 x) - \cos x = 1 \] \[ -2\cos^2 x - \cos x + 1 = 0 \] \[ 2u^2 + u - 1 = 0 \ \ \ \ \ ---- (*)\] where \( u = \cos x \) Now from the equation (*), \( u = 1/2 \ or -1 \). Now for what x is cos x = 1/2 or cos x = -1?

OpenStudy (jamesj):

If x is in the interval [0,2pi), cos x = -1 <=> x = pi cos x = 1/2 <=> pi/3, 5pi/3 Therefore the final answer is x=π/3,π,5π/3.

OpenStudy (phi):

@Jinnie, you should sketch out cos(x) between 0 and 2pi, and then mark all the places it takes on the above values. This way you won't miss any of the solutions

OpenStudy (anonymous):

oh wow james i had a near exact answer yesterday but i had 3pi/4 somewhere in there

OpenStudy (anonymous):

Find all solutions of the equation sec^(2)x−2=0. The answer is A+Bkπ where k is any integer and 0<A<π/2, Find A and B can you help me find just B thats all i need

OpenStudy (jamesj):

\[ \sec^2 x - 2 = 0 \iff \cos x = \pm 1/\sqrt{2}\] Now solve these seperately, cos x = 1/sqrt(2) <=> x = ...what? cos x = -1/sqrt(2) <=> x = ...what?

OpenStudy (anonymous):

this is what i ended up with when i tried to do this sec^(2) x−2=0 1/ cos^2 x = 2 2cos^2 x = 1 cos^2 x = 1/2 cos x = (+/-)1/√2 x = π/4 x= 3π/4 x=5π/4 x =7π/4 x =π/4 + 2kπ x= 3π/4 + 2kπ x=5π/4 + 2kπ x = 7π/4 + 2kπ

OpenStudy (phi):

The only thing I don't understand is how you can express all those solutions with A between 0 and pi/2??

OpenStudy (jamesj):

\[ x = \pi/4 + k\pi/2 \]